Find the equation in standard form of an ellipse with center at (0,0) minor axis of length 18, and foci at (0,-12) and (0,12).
a. (x^2/225)+(y^2/81)=1
b. (x^2/144)+(y^2/81)=1
c. (x^2/81)+(y^2/225)=1
d. (x^2/9)+(y^2/15)=1
e. (x^2/15)+(y^2/9)=1
Answer by lwsshak3(11628) 
 Find the equation in standard form of an ellipse with center at (0,0) minor axis of length 18, and foci at (0,-12) and (0,12).
a. (x^2/225)+(y^2/81)=1
b. (x^2/144)+(y^2/81)=1
c. (x^2/81)+(y^2/225)=1
d. (x^2/9)+(y^2/15)=1
e. (x^2/15)+(y^2/9)=1
**
As noted from foci, where x doesn't change but y does, given ellipse has a vertical major axis with standard form: (x-h)^2/b^2+(y-k)^2/a^2=1, (a>b)
2b=minor axis=18
b=9
b^2=81
..
c=12
c^2=a^2-b^2
144=a^2-81
a^2=144+81=225
a=√225=15
..
Equation:
x^2/81+y^2/225=1
c. is the correct ans.
| |
|
| |